The current through the 3 ohm resistor is 1A what
Answer #1

in the question two volts are given so the change in voltage is 9V-6V = 3V

according to ohm's law V=IR

So, 3V = i x 3 ohm

i = 1A

similarily, now replace 6V battery with 10V then the voltage difference is 10V-9V = 1V

1V = i X 3 ohm

i = 1/3 amp